OpenStudy (anonymous):

Let a and b denote the roots of the equation x^2 - 10x + 6 =0 let sn whre n a positive integer stands for a^n + b^n s(n+2)=? s4=? s4-s2s3=?

OpenStudy (anonymous):

if v find the roots of d equation then the question will b a cake walk? so let us focus on finding d roots first

OpenStudy (anonymous):

or wait s4= a^4+b^4=(a^2+b^2)^2-2(ab)^2 =((a+b)^2-2ab)^2-2(ab)^2 now sum of roots(a+b)= 10 nd product ab=6, put these values nd u will find ur answer

OpenStudy (anonymous):

did nt understand lol

OpenStudy (anonymous):

the answer for s4 = 7672......

OpenStudy (anonymous):

see this was kind of transformation, so u need to understand it , it is in simplest language, what i did is turned the function into a function of (a+b) nd ab, coz we know these values

OpenStudy (anonymous):

so@ @dg123 by solving that i wont get 7672

OpenStudy (anonymous):

u know what are the values of a+b and ab?

OpenStudy (anonymous):

a^4 + b^4 = (a^2)^2 + (b^2)^2 =(a^2)^2 + (b^2)^2 + 2a^2 b^2 -2a^2 b^2 =(a^2 -b^2)^2 + 2(a^2

OpenStudy (anonymous):

@mukushla we can get that from the eq

OpenStudy (anonymous):

@shubhamsrg i want to know how this came s4= a^4+b^4=(a^2+b^2)^2-2(ab)^2 =((a+b)^2-2ab)^2-2(ab)^2

OpenStudy (anonymous):

aha i got it....

OpenStudy (anonymous):

then wat is s(n+2)=?

OpenStudy (anonymous):

\[a^{n+2}+b^{n+2}\]

OpenStudy (anonymous):

nothing to do with it

OpenStudy (anonymous):

ok then s4-s2s3=?

OpenStudy (anonymous):

just notice \[s2=a^2+b^2=(a+b)^2-2ab \\ and \ \ \ s3=a^3+b^3=(a+b)(a^2+b^2-ab)=(a+b)((a+b)^2-3ab)\]

OpenStudy (anonymous):

got it thxx